OpenStudy (anonymous):

what are the foci of the graph 4x^2-9y^2=36 a.(+-6,0) b.(0,+-6 c.(+-3.6,0) d.(0,3.6)

OpenStudy (amistre64):

4/36 = 1/9 ; -9/36 = -4 a = 3 and b = 4, the foci should be at 5 from center then

OpenStudy (amistre64):

3 and 2 sorry lol

OpenStudy (amistre64):

sqrt(9+4) = sqrt(13) for c

OpenStudy (amistre64):

im thinking it c